OpenStudy (anonymous):

What are the x and y intercepts of -8x + 4y = 32

OpenStudy (anonymous):

x - intercepts set y = 0 and solve for x via\[-8x=32\] \[x=\frac{32}{-8}=-4\] so x intercept is (-4,0)

OpenStudy (amistre64):

divide it all by 32 and simplify :)

OpenStudy (anonymous):

y - intercept put x = 0 and solve \[4y=32\] \[y=\frac{32}{4}=8\]

OpenStudy (anonymous):

y - intercept is (0,4)

OpenStudy (anonymous):

sensei would find working with fractions = "simplifying"

OpenStudy (amistre64):

\[-8x + 4y = 32\] \[\frac{-8x}{32} + \frac{4y}{32} = \frac{32}{32}\] \[\frac{-x}{4} + \frac{y}{8} = 1\] when x = 0; y = 8 when y = 0; x= -4 :)

OpenStudy (amistre64):

im assuming you typoed your y intercept :)

OpenStudy (anonymous):

Refer to the attached visual aid.

OpenStudy (anonymous):

yes a typo sorry. should be (0,8)